OpenStudy (anonymous):

Write an equation for the quadratic function whose graph contains the given vertex and point Vertex: (1,3), Point: (-4,-27) Please explain how you come to the answer!

OpenStudy (anonymous):

y=a(x=1)^2+3

OpenStudy (anonymous):

a(x+1)^2+3

OpenStudy (anonymous):

then put in the other points

OpenStudy (anonymous):

-27=a(-4+1)^2+3

OpenStudy (anonymous):

sooo y=2(x-1)^2+3

OpenStudy (anonymous):

u'll get the slope

OpenStudy (anonymous):

yea thats using the vertex

OpenStudy (anonymous):

then u substitue the other points of the graph

OpenStudy (anonymous):

to get the slope

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

n ones u get the slope u use the slope and the vertex only to get the equation

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[-27=a(-4+1)^2+3 -27=9a=3 -30=9a a=3\]

OpenStudy (anonymous):

then use the info u have y=3.3(x-1)^2+3

OpenStudy (anonymous):

thats the equation]

OpenStudy (anonymous):

how does that go to a y=ax^2+bx+c form? i want to know how you fit the things into this form.

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

just open the brakets up

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

n the exponets

OpenStudy (anonymous):

i got the ans y=3.3x^2-6.6x+6.3

OpenStudy (anonymous):

k bye